OpenStudy (maddieayeeee):

A cylindrical gasoline storage tank on Mr. Edwards' farm needs to be painted. The tank is 8 feet long and has a diameter of 4 feet. If one gallon of paint covers 350 square feet, how many cans of paint will Mr. Edwards need?

OpenStudy (yournightmare):

8*4=?^2

OpenStudy (yournightmare):

32^2

OpenStudy (maddieayeeee):

Divide?

OpenStudy (maddieayeeee):

1024

OpenStudy (yournightmare):

yeah that seems off.

OpenStudy (yournightmare):

hold on. i think it might be 2^2

OpenStudy (maddieayeeee):

4

OpenStudy (yournightmare):

no its 8. 2*2*2= 8

OpenStudy (maddieayeeee):

Are you sure that's right answer?

OpenStudy (yournightmare):

yes

OpenStudy (yournightmare):

to double check tho @Kainui

Directrix (directrix):

Lateral Area of Cylinder is this: (base circumference) * height = 2*π* 2 * 8= 32π Now, find the areas of the top and bottom.